OpenStudy (anonymous):

Find the general solution of x'=(2, 1, -5, -2)x+(-cos t, sin t). ( 2 and 1 on left, -5 and -2 on right. -cos t on top, sin t on bottom. how do I find the particular solution for the 2x1 matrix using undetermined coefficients?)

OpenStudy (loser66):

\[x'=\left[\begin{matrix}2&-5\\1&-2\end{matrix}\right]x +\left[\begin{matrix}-cost\\sintt\end{matrix}\right]\]

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (loser66):

@Idealist @SithsAndGiggles is here, we are fine

OpenStudy (anonymous):

Which method is easier to use for the particular solution in this case? Undetermined coefficients or variation of parameters?

OpenStudy (loser66):

hey, why I don't see any thing pop up? I saw Sith wrote something

OpenStudy (anonymous):

how to find the general solution for 2x1 matrix?

OpenStudy (anonymous):

Homogeneous solution: \[\begin{align*}\begin{vmatrix}2-\lambda&-5\\1&-2-\lambda\end{vmatrix}=-(4-\lambda^2)+5&=0\\ 4-\lambda^2&=5\\ \lambda^2&=-1\\ \lambda&=\pm i \end{align*}\] Finding the eigenvectors: \[\lambda_1=i:\\ \begin{pmatrix}2-i&-5\\1&-2-i\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}~~\Rightarrow~~\vec{\eta_1}=\begin{pmatrix}2+i\\1\end{pmatrix}\] So the second eigenvector must be \(\vec{\eta_2}=\begin{pmatrix}2-i\\1\end{pmatrix}\).

OpenStudy (anonymous):

I'll have to review how to write up the final general solution using the eigenvectors. In any case, I would suggest continuing with undetermined coefficients, but only because I don't remember how to use variation of parameters.

OpenStudy (anonymous):

can you show me how to find it using undetermined coefficients?

OpenStudy (anonymous):

but for the first eigenvector, I got a=i, b=2/5i+1/5 and for the second eigenvector, I got a=-i, b=-2/5i+1/5.

OpenStudy (anonymous):

For the homogeneous part, you should have \[\vec{x_c}=C_1\begin{pmatrix}2\cos t-\sin t\\\cos t\end{pmatrix}+C_2\begin{pmatrix}2\sin t+\cos t\\\sin t\end{pmatrix}\] I'll go over the details of finding the first eigenvector: \[\lambda_1=i:\\ \begin{pmatrix}2-i&-5\\1&-2-i\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] Multiply the top row by \(2+i\) and the bottom row by -5: \[\begin{pmatrix}(2-i)(2+i)&-5(2+i)\\-5&5(2+i)\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\\ \begin{pmatrix}5&-5(2+i)\\-5&5(2+i)\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\\ \begin{pmatrix}5&-5(2+i)\\0&0\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] So you have \[5\eta_1=5(2+i)\eta_2\\ \eta_1=(2+i)\eta_2\] So, \(\eta_2=1~~\Rightarrow~~\eta_1=2+i\).

OpenStudy (anonymous):

wait a minute.

OpenStudy (anonymous):

how did you get 0, 0 on the bottom?

OpenStudy (anonymous):

Sorry, that didn't answer your question... The step was adding the rows together. I thought you were talking about the zero vector's zeros.

OpenStudy (anonymous):

I get it now. how about the second eigenvector?

OpenStudy (anonymous):

The second eigenvector isn't important in the long run, but since you know one you also know the other. They come in conjugate pairs, so you don't have to show the same work for it.

OpenStudy (anonymous):

I got the same 2 eigenvectors, now what?

OpenStudy (anonymous):

Now for the U.C. part of the answer. I'm still not sure how to approach it, so you'll have to give me a moment.

OpenStudy (anonymous):

okay.